OpenStudy (anonymous):

Find 4 consecutive even integers where the product of the two smaller numbers is 56 less than the product of the two larger numbers.

OpenStudy (ipwnbunnies):

Let 'x' be the smallest of the integers. Then, the next 3 integers will be (x+2), (x+4), and (x+6). The product of the two smaller numbers is 56 LESS than the product of the 2 larger. \[x(x+2) + 56 = (x+4)(x+6)\] Does that seem right? Set it up as an equation.

OpenStudy (anonymous):

I wish I could tell you if that seemed right, but I have no idea... which is why I am here :D you might have to keep going step by step to help me understand because integers are my weakness. x_x

OpenStudy (ipwnbunnies):

An integer is a positive or negative "whole number". So, 3, 92, -235, and 0 are all integers.

OpenStudy (anonymous):

Okay. Got that part. Now what?

OpenStudy (ipwnbunnies):

Apparently, 0 may or may not be an integer. >_> But whatever. We don't have to deal with 0. So, we need 4 consecutive even integers. We'll let the smallest of the integers be x.

OpenStudy (ipwnbunnies):

And, the next consecutive even integer, you must add 2 to it. Like, 2, 4, 6, etc. You have to add 2. So, x is the smallest integer. The next smallest is (x+2). The next, (x+4). And finally, the biggest is (x+6).

OpenStudy (anonymous):

Okay I see the 4 integers but why are we adding 2 to each one?

OpenStudy (ipwnbunnies):

Because that's what makes them both even and consecutive. Another example. Let x = 10. (10 + 2) = 12 (10 + 4) = 14 (10 + 6) = 16 10, 12, 14, 16 are all even, consecutive integers.

OpenStudy (anonymous):

Oh, duh... -Facepalm- I'm so dumb -_-

OpenStudy (ipwnbunnies):

It's kewl. :3 Now, we move on to creating our equation for 'the product of the two smaller numbers is 56 less than the product of the two larger numbers.'

OpenStudy (ipwnbunnies):

'The product of the two smaller numbers' : x(x+2) 'the product of the two larger numbers' : (x+4)(x+6) Since the product of the smaller numbers is 56 less than the bigger product, we'll have to add 56 to the smaller product to make them equal. \[x(x+2) + 56 = (x+4)(x+6)\]

OpenStudy (anonymous):

Okay, I'm following.

OpenStudy (anonymous):

Do I solve that equation and it gives me the answer or is there more to it?

OpenStudy (ipwnbunnies):

Yes, solve the equation for x. It'll be a quadratic equation. Use the even integer for x. Then, you can find the consecutive, even integers. Remember, we already said that 'x' equals the smallest of the even integers.

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Oh, wut. You already told me what you did in chat. Yas, x = 4. Therefore, the next 3 consecutive integers are 6, 8 and 10. Check your work. \[(4)(6) + 56 = (8)(10) ?\]

OpenStudy (anonymous):

It does not equal that D;

OpenStudy (ipwnbunnies):

.-. I'm confused now. Your answer to the original question is correct, lol.